# Cog-TiPRO: Iterative Prompt Refinement with LLMs to Detect Cognitive Decline via Longitudinal Voice Assistant Commands

### v3

@@ I Introduction
Traditional diagnostics for MCI rely on clinical assessments, including neuroimaging, laboratory tests, and standardized neuropsychological examinations such as the Montreal Cognitive Assessment (MoCA) [4]. While effective for diagnosis, these methods have limitations: They require administration by trained healthcare professionals in clinical settings, making them labor-intensive and expensive for frequent monitoring. Since speech impairments such as increased pauses, word-finding difficulty, and language disfluency are among the earliest symptoms of cognitive decline, speech processing represents a promising approach for non-invasive and cost-effective MCI detection that could potentially be implemented in the everyday environment [5].
@@ III-A Data Collection Details
Participants were classified as MCI or HC based on MoCA scores ($\geq$ 26 for HC, << 26 for MCI). In the initial assessment, 8 were MCI and 7 were HC (labels). Ten participants had prior VAS experiences. All participants had average MDPQ-16 (Mobile Device Proficiency Questionnaire) [15] score above 4.5/5, indicating that they had adequate technical proficiency. The study combined two data collection elements:
@@ VII-B Ablation Study of Main Components
Figure 2: Prompt optimization progress: initial template (left), intermediate template (middle), and final optimized prompt (right). The bottom panel shows an example of error analysis feedback for prompt edits.
@@ IX Conclusion
In this paper, we demonstrate the viability of using VAS for the longitudinal data collection and analysis of voice commands to detect cognitive decline. By collecting daily interactions from 15 older adults over 18 months in their homes, we capture speech patterns that reflect cognitive functioning through our Cog-TiPRO framework. Cog-TiPRO leverages LLMs to extract linguistic features and aspects indicative of MCI from unstructured, short, and noisy voice commands. The integration of LLM-driven iterative prompt refinement, acoustic feature extraction, and temporal modeling enables more accurate detection of MCI. Using iTransformer, our approach achieves 73.80% accuracy and 72.67% F1-score, outperforming its baseline by 27.13%. The linguistic features extracted not only reveal patterns specific to everyday VAS interactions but also align with established clinical markers of MCI, demonstrating the potential of passive monitoring for the early detection of cognitive decline.

### v4

@@ I Introduction
Traditional diagnostics for MCI rely on clinical assessments, including neuroimaging, laboratory tests, and standardized neuropsychological examinations such as the Montreal Cognitive Assessment (MoCA) [4]. While effective for diagnosis, these methods have limitations: They require administration by trained healthcare professionals in clinical settings, making them labor-intensive and expensive for frequent monitoring. Since speech impairments such as increased pauses, word-finding difficulty, vocal tremors, and weaker coherence are among the earliest symptom markers of cognitive decline [5, 6, 7], speech processing represents a promising approach for non-invasive and cost-effective MCI detection that could potentially be implemented in the everyday environment.
@@ III-A Data Collection Details
Participants were classified as MCI or HC based on MoCA scores ($\geq$ 26 for HC, << 26 for MCI). In the initial assessment, 8 were MCI and 7 were HC (labels). Ten participants had prior VAS experiences. All participants had an average MDPQ-16 (Mobile Device Proficiency Questionnaire) [17] score above 4.5/5, indicating that they had adequate technical proficiency. The study combined two data collection elements:
@@ VII-B Ablation Study of Main Components
Figure 2: Prompt optimization progress: initial template by a human (left), intermediate template after the first iteration (middle), and final optimized prompt after three iterations (right). The bottom panel shows an example of error analysis feedback for prompt edits.
@@ IX Conclusion
In this paper, we demonstrate the viability of using VAS for the longitudinal data collection and analysis of voice commands to detect cognitive decline. By collecting daily interactions from 15 older adults over 18 months in their homes, we capture speech patterns that reflect cognitive functioning through our Cog-TiPRO framework. Cog-TiPRO leverages LLMs to extract linguistic features and aspects indicative of MCI from unstructured, short, and noisy voice commands. The integration of LLM-driven iterative prompt refinement, acoustic feature extraction, and temporal modeling enables more accurate detection of MCI. Using iTransformer, our approach achieves 73.80% accuracy and 72.67% F1-score, outperforming its baseline by 27.13%. The linguistic features extracted not only reveal patterns specific to everyday VAS interactions but also align with established markers of MCI, such as word-finding difficulties, disfluency, and weaker coherence. This finding demonstrates the potential of passive monitoring for the early detection of cognitive decline.